what is the subset {17,1,14,12}
number of proper subsets

Answer by stanbon(75887)

{17,1,14,12}
---
Total # of subsets = 2^4
---
Total # of proper subsets = 2^4 -1 = 15
===========================================
Cheers,
Stan H.